OpenStudy (anonymous):

Using a directrix of y = 5 and a focus of (4, 1), what quadratic function is created?

OpenStudy (anonymous):

f(x) = 1/4 (x − 4)2 − 3 f(x) = 1/8 (x + 4)2 − 3 f(x) = −1/8 (x − 4)2 + 3 f(x) = −1/4 (x + 4)2 − 3

OpenStudy (anonymous):

I think it's A or C

OpenStudy (anonymous):

@amity hey again :)

OpenStudy (amity):

Yes i am reading your problem

OpenStudy (amity):

Yes i am reading your problem

OpenStudy (anonymous):

thanks

OpenStudy (amity):

It is a parabola not a ellipse and hyperbola as single directrix

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it's C?

OpenStudy (amity):

It is a parabola not a ellipse and hyperbola as single directrix

OpenStudy (anonymous):

um..I know that

OpenStudy (anonymous):

I'm just confused. so it's A instead?

OpenStudy (amity):

Answer is c

OpenStudy (amity):

Answer is c

OpenStudy (anonymous):

thank you xx

OpenStudy (amity):

Take any point (x,y) and its distance from focus and directrix equate

OpenStudy (amity):

Take any point (x,y) and its distance from focus and directrix equate

OpenStudy (amity):

Welcome yyyy

OpenStudy (amity):

Welcome yyyy